# Cryogenic Thermal Modeling of Microwave High Density Signaling

### v1

@@ Abstract
Superconducting quantum computers require microwave control lines running from room temperature to the mixing chamber of a dilution refrigerator. Adding more lines without preliminary thermal modeling to make predictions risks overwhelming the cooling power at each thermal stage. In this paper, we investigate the thermal load of SC-086/50-SCN-CN semi-rigid coaxial cable, which is commonly used for the control and readout lines of a superconducting quantum computer, as we increase the number of lines to a quantum processor. We investigate the makeup of the coaxial cables, verify the materials and dimensions, and experimentally measure the total thermal conductivity of a single cable as a function of the temperature from cryogenic to room temperature values. We also measure the cryogenic DC electrical resistance of the inner conductor as a function of temperature, allowing for the calculation of active thermal loads due to Ohmic heating. Fitting this data produces a numerical thermal conductivity function used to calculate the static heat loads due to thermal transfer within the wires resulting from a temperature gradient. The resistivity data is used to calculate active heat loads, and we use these fits in a cryogenic model of a superconducting quantum processor in a typical Bluefors XLD1000-SL dilution refrigerator, investigating how the thermal load increases with processor sizes ranging from 100 to 225 qubits. We conclude that the theoretical upper limit of the described architecture is approximately 200 qubits. However, including an engineering margin in the cooling power and the available space for microwave readout circuitry at the mixing chamber, the practical limit will be approximately 140 qubits.
@@ I Introduction
Superconducting quantum computers use qubits with energy gaps between the ground and the first excited states in the microwave regime. Frequencies in this range allow for qubit manipulation and readout using microwave components in the low-photon number regime but also require the devices to be operated at temperatures of order 20 mK so that the thermal population of the first excited state is negligible. Commercially available dilution refrigerators achieve these temperatures, but the system must be engineered such that the total heat load coming from the outside or generated within the cryostat is lower than the cooling power available from the refrigeration system. The qubit gate operation and readout are controlled by microwave pulses generated by room-temperature electronics outside the cryostat. These pulses travel down through signaling cables that are attached at the different temperature stages of the dilution refrigerator. The cables create a static heat load due to thermal conductivity and temperature differences between stages. Any cable that carries a current also has an active load due to Ohmic heating due to the cable resistivity and heat loss in attenuators used to thermalize the cable at each stage Clarke and Wilhelm (2008).
@@
Scaling the number of qubits to the 100-200 range requires an increase in the number of signal lines; however, there is a limit to the number of cables that can be accommodated for any one particular cable type. This limiting factor is usually either the thermal budget or the available space Krinner et al. (2019). More compact and dense microwave cable types are being actively investigated Deshpande et al. (2019), and other potential solutions to scaling the number of control signals have been reported, such as using cryo-CMOS circuits either for RF switching Potočnik et al. (2021) or arbitrary waveform generation within the cryostat Pellerano et al. (2022), or using optical fibers and a transducer to create RF pulses from optical signals Lecocq et al. (2021).
@@
This work extends the work of Krinner et al. Krinner et al. (2019) in modeling the thermal loads resulting from the delivery of signals to superconducting quantum processors. Here, we examine a commonly used small-format coaxial cable: the SC-086/50-SCN-CN semi-rigid coaxial cable used in the Bluefors High-Density Wiring (HDW). These cables have the same internal dimensions as the UT-085 cables considered by Krinner et al. (2019), with a cupronickel outer and silver-coated cupronickel inner conductor with a PTFE dielectric. Assemblies of these coaxial lines are used in the Bluefors XLD1000-SL range of dilution refrigerators to allow simplified wiring configuration changes.
@@
In this paper, we determine the material composition and cross-sectional dimensions of the different parts of the cable. The thermal conductivity of the various parts was then experimentally measured as a function of temperature, and we then developed a thermal model of the coaxial cable static heat load across all thermal stages spanned by the cable. In addition, we measured the electrical resistivity of the signal line as a function of temperature to estimate the active loads due to Ohmic heating for a single cable. This active load is determined by the system configuration and architecture, where the current in the signal line and the attenuators dissipate heat. Finally, we combine the static and active heat load models to simulate a complete superconducting quantum computer ranging in size from 100 to 225 qubits.
@@ II Overview of Thermal Modeling
Current designs of dilution refrigerators are able to cool the sample space to under 20 mK without cryoliquids using a He3/He4 dilution unit that is precooled by a two-stage pulse tube refrigerator Uhlig (2002, 2008). The cables within a dilution refrigerator are thermally anchored at several stages within the fridge whose temperatures are set by the active cooling systems. The nominal stage temperatures and estimated cooling powers for the Bluefors XLD1000-SL system, along with the cable lengths, are shown in Table 1. Bluefors specifies the cooling power and base temperature only at the MXC stage Bluefors (2025), and the numbers for higher temperature stages are estimates based on our own experience of running these systems and are indicated by an asterisk in Table 1. These cooling powers are those available to cool experimental loads added to the bare refrigerator. The dilution fridge is a coupled dynamic system, and the estimated cooling power of each fridge stage varies according to how the system is operated and optioned, depending on the specifications of the installed pulse tube coolers and the dilution units. Pulse tube technology has improved substantially over the past years, and the available cooling power is expected to increase. For this paper, we assume the system has two Cryomech PT420 pulse tubes, the standard configuration of the Bluefors XLD1000-SL at the time of writing. Bluefors currently has the option to upgrade these to the Cryomech PT425 or PT310 pulse tube configurations for greater cooling power and additional options to have three pulse tubes at the expense of reducing the number of side-loading ports for the wiring Bluefors (2025). The fridge has multiple wiring options, and here we focus on the SC-086/50-SCN-CN coax cable used in the Bluefors High-Density Wiring (HDW) option.
@@
The calculated static and active heat loads for each model are summed to provide a total input heat load at each stage. The final heat load at each fridge stage is determined by the difference between the incoming heat load from higher temperature stages, the Ohmic heating, and the outgoing load to the next lower temperature stage. This final load is then compared to the estimated cooling power of each fridge stage shown in Table 1.
@@
To model a complete superconducting processor, we take a common design that uses flux tunable transmons as the basic qubit design, with tunable couplers between pairs of qubits that are designed for use in two-qubit gate operations Yan et al. (2018). The control and readout lines for these processors can be split into four different categories:
@@
The different functional groups of the microwave lines will each have a different attenuator configuration to ensure the thermal noise on the signal line at the processor is at the fridge’s base temperature. All of the lines provide a static heat load because they run from room temperature to the MXC, while the active Ohmic load depends on the current in the line and the attenuator configuration.
The qubit XY lines carry infrequent microwave pulses for single-qubit manipulation, while the flux bias line has a constant current to set the qubit frequency. The tunable couplers have a similar flux bias line with a constant current to set the operating point so that the two qubits do not interact, which is changed to enable two-qubit gate interactions. Each qubit is coupled to a resonator for readout, and the readout circuit has a read-in line coupled to several resonators and continues on to a readout line with isolators and amplifiers. A traveling wave parametric amplifier (TWPA) and a low noise semiconductor amplifier are frequently used to boost the output signal in the readout. The TWPA requires a separate microwave line to provide a microwave pump signal, and a semiconductor amplifier will have twisted-pair dc wiring and produce heat that must be accounted for.
@@
Table 1: Estimated temperatures and cooling power at each refrigerator temperature stage of a Bluefors XLD1000-SL system for a bare fridge. The acronyms “CP” and “MXC” respectively refer to the cold plate and mixing chamber stages of the dilution fridge. Cable lengths refer to the cable run from the higher temperature stage. “Below MXC” refers to the cables run from the MXC to the package holding the quantum processor. The Below MXC cables are not considered in the static heat load as they are isothermal at the mixing chamber temperature. They do, however, contribute an active load. Bluefors specifies the cooling power and base temperature only at the MXC stage Bluefors (2025); the numbers for higher temperature stages are estimates based on our own experience of running these systems, and they are indicated by an asterisk.
@@ III.1 Overview of Static Heat Loads
Coaxial cables carry heat from the hotter fridge stages, and each material component has a different thermal conductivity. We model the inner and outer conductors and the dielectric layer of the coaxial cable as separate parallel thermal paths, which we then sum to produce a total static heat load. The model requires the dimensions of the material components, the fixed temperatures at each end of the cable, and each material’s thermal conductivity as a function of temperature.
@@
While comprehensive thermal conductivity functions exist for various grades of copper and aluminum, models for certain cupronickel alloys are not generally available from cryogenic to room temperatures. In this study, we measure the thermal conductivity of the outer and inner conductors of SC-086/50-SCN-CN coaxial cable samples (Coax Co., Ltd.) as a function of temperature from cryogenic to room temperature. For the thermal conductivity of the PTFE dielectric layer, we use the published Teflon $k(T)$ function from the NIST Cryogenics Index of Material Properties database Marquardt et al. (2002); Bradley, P.E. et al. (2006). While PTFE and Teflon are both ways to refer to Polytetrafluoroethylene, we will refer to this material as “PTFE”.
@@
The static model only considers heat transport by thermal conduction. While infrared radiation is present, the cryogenic temperatures and the low emissivity of the outer conductor of SC-086/50-SCN-CN coaxial cable ensure it is insignificant compared to the thermal conductivity Krinner et al. (2019). The only time the correction due to infrared radiation is accounted for is in the experimental measurement of thermal conductivity where the Quantum Design software incorporates an estimate of the radiative losses into the measurement, and thus we must ensure that the emissivity is set to a realistic value.
III.2 Composition and Construction of SC-086/50-SCN-CN Coaxial Cables
@@
Figure 1: (a) Depiction of the layers (not to scale) of an SC-086/50-SCN-CN coaxial cable (Coax Co., Ltd.) with materials and dimensions taken from the manufacturers’ website CoaxCo . (b) The SEM cross-section of an SC-086/50-SCN-CN coaxial cable sample from Coax Co., Ltd. has measured dimensions that agree with the reported dimensions in 1(a). Measured dimensions and scale are enlarged for readability. (c) A higher magnification image of the inner conductor of an SC-086/50-SCN-CN coax cable sample. The measured silver coating is approximately 3 $\mu m$ thick, which agrees with the nominal thickness reported in CoaxCo . Measured dimensions and scale are enlarged for readability. (d) An enlarged portion of 1(c) to make the dimension of the silver coating readable. A larger decimal point was digitally added to the image for clarity.
The SC-086/50-SCN-CN cable (Coax Co., Ltd.) has a C7150 cupronickel outer conductor, a PTFE dielectric, and a silver-coated (nominal thickness= 3 - 4 $\mu$$m$) C7150 inner conductor. The material specifications and dimensions are detailed in Table 2 and are taken from CoaxCo .
@@
Fig. 1(a) shows a diagram of the construction of the cable along with nominal dimensions, and Fig. 1(b) and 1(c) shows cross-sectional scanning electron microscope pictures of a cable sample confirming the dimensions, including the thickness of the silver plating on the signal line. The composition of the cupronickel inner and outer conductors, as well as the dielectric, were measured using Energy Dispersive X-ray (EDX) analysis to determine the atomic ratios of materials. Table 3 gives the measured atomic ratios of copper to nickel for the cupronickel alloy in both conductors, which is approximately 62.5:37.5, and the measured atomic ratio of carbon to fluorine for the dielectric, which is close to the 1:2 ratio expected from the (C2F4)n molecular structure of PTFE.
@@
Table 2: Dimensional and Material Specifications of SC-086/50-SCN-CN Coaxial Cable CoaxCo . The cross-sectional areas of the material in each layer are calculated from the diameters of the different components, subtracting the area of the inner layers for the outer and dielectric layers.
@@
Figure 2: The cupronickel outer conductor of an SC-086/50-SCN-CN coaxial cable mounted on the PPMS TTO measurement system, showing the sample heater at the left used by the system to produce a time dependent thermal gradient across the sample, followed by two thermometers. The sample is thermally anchored at the right-hand end.
III.3 Experimental Measurement of Thermal Conductivity
A sample of SC-086/50-SCN-CN coaxial cable was deconstructed into its three constituent components: the C7150 cupronickel outer conductor, the PTFE dielectric, and the silver-coated C7150 cupronickel inner conductor.

### v3

@@ Abstract
Superconducting quantum computers require microwave control lines running from room temperature to the mixing chamber of a dilution refrigerator. Adding more lines without preliminary thermal modeling to make predictions risks overwhelming the cooling power at each thermal stage. In this paper, we investigate the thermal load of SC-086/50-SCN-CN semi-rigid coaxial cable, which is commonly used for the control and readout lines of a superconducting quantum computer, as we increase the number of lines to a quantum processor. We investigate the makeup of the coaxial cables, verify the materials and dimensions, and experimentally measure the total thermal conductivity of a single cable as a function of the temperature from cryogenic to room temperature values. We also measure the cryogenic DC electrical resistance of the inner conductor as a function of temperature, allowing for the calculation of active thermal loads due to Ohmic heating. Fitting this data produces a numerical thermal conductivity function used to calculate the static heat loads due to thermal transfer within the wires resulting from a temperature gradient. The resistivity data is used to calculate active heat loads, and we use these fits in a cryogenic model of a superconducting quantum processor in a typical Bluefors XLD1000-SL dilution refrigerator, investigating how the thermal load increases with processor sizes ranging from 100 to 225 qubits. We conclude that the theoretical upper limit of the described architecture is approximately 200 qubits. However, including an engineering margin in the cooling power and the available space for microwave readout circuitry at the mixing chamber, the practical limit is approximately 140 qubits.
@@ I Introduction
Superconducting quantum computers use qubits with energy gaps between the ground and the first excited states in the microwave regime. Frequencies in this range allow for qubit manipulation and readout using microwave components in the low-photon number regime, but also require the devices to be operated at temperatures of order 20 mK so that the thermal population of the first excited state is negligible. Commercially available dilution refrigerators achieve these temperatures, but the system must be engineered such that the total heat load coming from the outside or generated within the cryostat is lower than the cooling power available from the refrigeration system. The qubit gate operation and readout are controlled by microwave pulses generated by room-temperature electronics outside the cryostat. These pulses travel down through signaling cables that are attached at the different temperature stages of the dilution refrigerator. The cables create a static heat load due to thermal conductivity and temperature differences between stages. Any cable that carries a current also has an active load due to Ohmic heating, due to the cable resistivity, and heat loss in attenuators used to thermalize the cable at each stage clarke2008.
@@
Scaling the number of qubits requires an increase in the number of signal lines; however, there is a limit to the number of cables that can be accommodated for any one particular cable type. This limiting factor is usually either the thermal budget or the available space krinner2019. More compact and dense microwave cable types are being actively investigated deshpande2019, and other potential solutions to scaling the number of control signals have been reported, such as using optical fibers and a transducer to create RF pulses from optical signals lecocq2021, or using cryogenic complementary metal oxide semiconductor (cryo-CMOS) circuits either for RF switching potocnik2021 or arbitrary waveform generation within the cryostat pellerano2022. Still more solutions have been explored in recent years, including those that employ low-temperature superconducting digital circuits for control, such as superconductive rapid single flux quantum (RSFQ) and adiabatic quantum flux parametron (AQFP) circuits, both of which promise high energy efficiency krylov2024single. Theoretical optimization-based techniques, such as those described in Zhuldassov et al. Zhuldassov2023, can also be used, which identify the optimal operating temperatures for circuit components in a cryogenic computing system.
@@
While all of these techniques show promise, an understanding of the thermal properties of the materials that comprise a signaling cable is necessary, and thermal conductivity and resistivity data allows for the strategic planning of wiring configurations such that the total power dissipation is minimized. This work extends the analysis and data presented in Krinner et al. krinner2019 in modeling the thermal loads resulting from the delivery of signals to superconducting quantum processors. We offer additional insight by discussing the viability of processor scaling to 100-225 qubits and by focusing our study on a widely used and commercially available small-format coaxial cable for which detailed datasets did not previously exist. The particular cable is the SC-086/50-SCN-CN semi-rigid coaxial cable used in the Bluefors High-Density Wiring (HDW). These cables have the same internal dimensions as the UT-085 cables considered by krinner2019, with a cupronickel outer and silver-coated cupronickel inner conductor with a Polytetrafluoroethylene (PTFE) dielectric. Assemblies of these coaxial lines are used in the Bluefors XLD1000-SL range of dilution refrigerators to allow simplified wiring configuration changes.
@@
In this work, Section II describes a general overview of how we approach the thermal modeling, including the cooling powers available at different fridge stages and the details of the different control lines used to control a superconducting processor. Section III.1 then details the material composition and cross-sectional dimensions of the inner signal line, dielectric, and outer ground of the SC-086/50-SCN-CN semi-rigid coaxial cable. In Section III.2, we report the measured thermal conductivities of the inner signal and outer ground sections as a function of temperature, along with fits to the data. In Section III.4, we use the measured thermal conductivity fits with data for the thermal conductivity of the dielectric layer published by NIST marquardt2002; bradleyp.e.2006, to develop a thermal model of the static heat load of the coaxial cable across all thermal stages spanned by the cable. We then consider the active loads due to Ohmic heating, beginning with Section IV.1, in which we report the measured resistivity of the signal conductor of an SC-086/50-SCN-CN cable as a function of temperature along with fits to the data. Sections IV.2 and IV.3 describe the active load thermal model and calculations, respectively that use the resistivity data. The active load is determined by the system configuration and architecture, where the current in the signal line and the attenuators dissipate heat. Finally, in Section V, we combine the static and active heat load models to simulate a complete superconducting quantum computer ranging in size from 100 to 225 qubits.
@@ II Overview of Thermal Modeling
Current designs of dilution refrigerators are able to cool the sample space to under 20 mK without cryoliquids using a He3/He4 dilution unit that is precooled by a two-stage pulse tube refrigerator uhlig2002; uhlig2008. The cables within a dilution refrigerator are thermally anchored at several stages within the fridge, whose temperatures are set by the active cooling systems. The nominal stage temperatures and estimated cooling powers for the Bluefors XLD1000-SL system, along with the cable lengths, are shown in Table 1. Bluefors specifies the cooling power and base temperature only at the Mixing Chamber (MXC) stage Bluefors2025, and the numbers for higher temperature stages are estimates based on “heat maps” we have measured, where we position heaters on the various stages of a bare fridge and record the variation of stage temperatures in response to varied heater power. These cooling powers are indicated by an asterisk in Table 1, and they are those available to cool experimental loads added to the bare refrigerator. The cooling powers in Table 1 are close to those reported by Krinner krinner2019 for the smaller Bluefors LD400 system with suitable scaling for the more powerful dilution unit. The main differences are at the 4K stage, where Krinner reports 1.5 W of cooling power at 4.2 K, whereas we use a figure of 0.7 W at 3.5 K. There are notable differences in the pulse tubes and fridge hardware here, but the lower cooling power is mainly attributable to the lower operating point temperature.
The dilution fridge is a coupled dynamic system, and the estimated cooling power of each fridge stage varies according to how the system is operated and optioned, depending on the specifications of the installed pulse tube coolers and the dilution units. Pulse tube technology has improved substantially over the past years, and the available cooling power is expected to increase. For this paper, we assume the system has two Cryomech PT420 pulse tubes, which were the standard configuration of the Bluefors XLD1000-SL at the time of model development. However, it should be noted that, as of this writing, the current default configuration uses two Cryomech PT425 pulse tubes with an option to replace one of the Cryomech PT425 pulse tubes with a Cryomech PT310 pulse tube Bluefors2025. Another option is the use of three pulse tubes at the expense of reducing the number of side-loading ports for the wiring Bluefors2025. The fridge has multiple wiring options, and here we focus on the SC-086/50-SCN-CN coax cable used in the Bluefors High-Density Wiring (HDW) option.
@@
In this paper, we calculate and sum the static and active heat loads for each model to provide a total input heat load at each stage. We then calculate the final heat load at each fridge stage, taking into account the difference between the incoming heat load from higher temperature stages, the Ohmic heating, and the outgoing load to the next lower temperature stage. We compare this final load to the estimated cooling power of each fridge stage shown in Table 1.
@@
To model a complete superconducting processor, we take a common design that uses flux-tunable transmons as the basic qubit design, with tunable couplers between pairs of qubits that are designed for use in two-qubit gate operations yan2018. The control and readout lines for these processors can be split into four different categories:
@@
The different functional groups of the microwave lines each have a different attenuator configuration to ensure the thermal noise on the signal line at the processor is at the base temperature of the fridge. All of the lines provide a static heat load because they run from room temperature to the MXC, while the active Ohmic load depends on the current in the line and the attenuator configuration.
In this model design, the qubit XY lines carry infrequent microwave pulses for single-qubit manipulation, while the flux bias line has a constant current to set the qubit frequency. The tunable couplers have a similar flux bias line with a constant current to set the operating point so that the two qubits do not interact, which is changed to enable two-qubit gate interactions. Each qubit is coupled to a resonator for readout, and the readout circuit has a read-in line coupled to several resonators and continues on to a readout line with isolators and amplifiers. A Traveling Wave Parametric Amplifier (TWPA) and a low noise semiconductor amplifier are frequently used to boost the output signal in the readout. The TWPA requires a separate microwave line to provide a microwave pump signal, and a semiconductor amplifier has twisted-pair DC wiring and produces heat that must be accounted for.
@@
Table 1: Estimated temperatures and cooling power at each refrigerator temperature stage of a Bluefors XLD1000-SL system for a bare fridge. Cable lengths refer to the cable run from the higher temperature stage. “Below MXC” refers to the cables run from the MXC to the package holding the quantum processor. The Below MXC cables are not considered in the static heat load as they are isothermal at the mixing chamber temperature. They do, however, contribute an active load. Bluefors specifies the cooling power and base temperature only at the MXC stage Bluefors2025; the numbers for higher temperature stages are estimates based on our own tests of stage temperature variation in response to varied heating power, and they are indicated by an asterisk.
@@ III Static Heat Loads
Coaxial cables carry heat from the hotter fridge stages, and each material component has a different thermal conductivity. We model the inner and outer conductors and the dielectric layer of the coaxial cable as separate parallel thermal paths, which we then sum to produce a total static heat load. The model requires the dimensions of the material components, the fixed temperatures at each end of the cable, and the thermal conductivity of each material as a function of temperature.
@@
While comprehensive thermal conductivity functions exist for various grades of copper and aluminum, models for certain cupronickel alloys are not generally available from cryogenic to room temperatures. In this study, we measure the thermal conductivity of the outer and inner conductors of SC-086/50-SCN-CN coaxial cable samples (Coax Co., Ltd.) as a function of temperature from cryogenic to room temperature. For the thermal conductivity of the PTFE dielectric layer, we use the published Teflon $k(T)$ function from the NIST Cryogenics Index of Material Properties database marquardt2002; bradleyp.e.2006. While PTFE and Teflon are both ways to refer to Polytetrafluoroethylene, in the present paper, we refer to this material as “PTFE”.
@@
The static model only considers heat transport by thermal conduction. While infrared radiation is present, the cryogenic temperatures and the low emissivity of the outer conductor of SC-086/50-SCN-CN coaxial cable ensure it is insignificant compared to the thermal conductivity krinner2019. The only time we account for the correction due to infrared radiation is in the experimental measurement of thermal conductivity, where the Quantum Design software incorporates an estimate of the radiative losses into the measurement, and thus, we must ensure that we set the emissivity to a realistic value.
III.1 Composition and Construction of SC-086/50-SCN-CN Coaxial Cables
@@
Figure 1: (a) Depiction of the layers (not to scale) of an SC-086/50-SCN-CN coaxial cable (Coax Co., Ltd.) with materials and dimensions taken from the manufacturer’s website sc-08650. (b) The SEM cross-section of an SC-086/50-SCN-CN coaxial cable sample from Coax Co., Ltd. has measured dimensions that agree with the reported dimensions in 1(a). Measured dimensions and scale are enlarged for readability. (c) A higher magnification image of the inner conductor of an SC-086/50-SCN-CN coax cable sample. The measured silver coating is approximately 3 $\mu m$ thick, which agrees with the nominal thickness coaxcocomm. Measured dimensions and scale are enlarged for readability. (d) An enlarged portion of 1(c) to make the dimension of the silver coating readable. A larger decimal point was digitally added to the image for clarity.
[…]
@@
A diagram of the construction of the cable along with nominal dimensions is shown in Fig. 1(a). Cross-sectional scanning electron microscope pictures of a cable sample are shown in Figs. 1(b) and 1(c), which confirm the cable dimensions, including the thickness of the silver plating on the signal line. The composition of the cupronickel inner and outer conductors, as well as the dielectric, were measured using Energy Dispersive X-ray (EDX) analysis to determine the atomic ratios of materials. The measured atomic ratios of copper to nickel for the cupronickel alloy in both conductors, which is approximately 62.5:37.5, and the measured atomic ratio of carbon to fluorine for the dielectric, which is close to the 1:2 ratio expected from the (C2F4)n molecular structure of PTFE, are listed in Table 3.
@@
Table 2: Dimensional and Material Specifications of SC-086/50-SCN-CN Coaxial Cable sc-08650. The cross-sectional areas of the material in each layer are calculated from the diameters of the different components, subtracting the area of the inner layers for the outer and dielectric layers.
@@
Figure 2: The cupronickel outer conductor of an SC-086/50-SCN-CN coaxial cable mounted on the PPMS TTO measurement system. From left to right, we see the sample heater, which produces a time-dependent thermal gradient across the sample, two thermometers, and then thermal anchoring of the sample to the right-hand end.
III.2 Experimental Measurement of Thermal Conductivity